# Quasicrystals as an intermediate form of matter between crystalline and amorphous solids

### v1

@@ II.3 Quasicrystal structure and dynamics
III Results
III.1 Thermodynamic properties and structure
III.1.1 Density
During the rapid cooling process from the high-temperature liquid phase, the density displays two distinct transitions, as illustrated in Fig. 5. The first transition occurs at approximately $T_{LH}=1.78$, marking the shift from the liquid phase to the orientationally correlated hexatic phase. Subsequently, the second transition takes place as the system undergoes a transformation from the hexatic phase to a dodecagonal quasicrystal phase, occurring at around $T=0.87$. It is worth noting that the transition from the hexatic to the quasicrystal phase is characterized by a positive thermal expansion. In contrast, during the transition from the liquid phase to the hexatic phase, there is no observable expansion or compression, as depicted in Fig. 5.
III.1.2 Heat capacity
We calculate the heat capacity using the ”direct” method of taking the derivative of the total energy with respect to temperature in cooling paths. Although this method, while applicable in non-equilibrium heating/cooling paths, may large uncertainties in estimating thermodynamic properties in equilibrium [86]. It is sufficient for calculating the relative value of different phases and representing phase transitions. In a manner similar to the density, the heat capacity as a function of temperature (Fig. 6) reveals two well-defined thermodynamic transitions. These transition temperatures ($T_{LH}=1.78$ and $T_{HQ}=0.87$) are consistent with previous findings. The two transitions seem sharp, but it is not possible to deduce the transition order due to the limited system size as above mentioned.
@@ III.1.2 Heat capacity
III.1.3 Radial distribution function
The RDF is a widely utilized tool for elucidating the structural characteristics of ordered and disordered materials. In Fig. 7, we present the RDF for the liquid phase at $T=1.95$, the hexatic phase at $T=1.2$, and the dodecagonal quasicrystal at $T=0.3$. The RDF of the quasicrystal exhibits a distinctive behavior that positions it between the characteristics of crystalline and disordered materials. This observation supports the notion of an intermediate state of matter, as alluded to in the title and introduction of our paper.
In addition to the well-established RDF profiles in amorphous and hexatic phases, the quasicrystal exhibits unique features. In the short range, the RDF of the quasicrystal reveals distinct peaks, which can be attributed to the quasi-periodic tiling. The first peak emerges at approximately $r=1.0$, while the second peak appears around $r=1.89$. These values are readily explicable as they correspond to the positions of potential energy wells (stable states) within the LJG-type potential (refer to Fig. 2(a)). Conversely, at longer distances, the RDF of the quasicrystal displays typical characteristics of disordered materials. It gradually converges towards a constant value with non-vanishing fluctuations, reflecting the absence of long-range translational symmetry, akin to amorphous systems, while still retaining certain translational symmetry.
@@ III.1.3 Radial distribution function
III.1.4 Static structure factor
@@
As depicted in Fig. 8, the static structure factors of the liquid, hexatic, and quasicrystal phases all prominently feature a main peak around $k^{*}=2\pi/a$, where $a$ represents the nearest-neighbor distance in real space. A pronounced peak in the structure factor at $k^{*}$ serves as a indication of local correlations in the particle positions in the condensed material. This peak corresponds to a wave vector that generates a long-range oscillatory behavior with a periodicity of $a$ in real space [88]. Moreover, the height of the first peak in the static structure factor is a critical parameter for predicting the liquid-solid phase transition. According to the Hansen-Verlet criterion in crystalline materials the freezing transition occurs when the value of the first peak in $S(k)$ reaches a specific constant value [89]. In the liquid phase, the first peak is notably lower than those in the hexatic and quasicrystal phases, underscoring the lack of long-range order and the inherent disorder in the system [90].
@@
Figure 8: Static structure factor of the (a) liquid phase at $T=1.95$, (b) hexatic phase at $T=1.2$ and (c) dodecagonal quasicrystal at $T=0.3$.
We have also generated the intensity maps of the static structure factor using Eq. (5), which provide a comprehensive representation of the 2D structures included in the phase diagram. These intensity maps resemble the non-energy-resolved diffraction patterns typically obtained in diffraction experiments to determine crystal structure and symmetry. The diffraction pattern of the disordered phase (Fig. 9(a)) consists of diffuse rings of light, resulting from the lack of regular arrangement in ordered lattices. In contrast, the diffraction pattern of the hexatic phase exhibits both weak polyhedra and bright reflections with 6-fold and 12-fold symmetries (Fig. 9(b)), clearly indicating the hexatic phase intermediate nature between the liquid and solid phases. The diffraction pattern of the quasicrystal (Fig. 9(b)) is more intricate, displaying a 12-fold symmetry and well-defined reflections arranged in a quasi-periodic 2D pattern. These diffraction patterns serve to reinforce the credibility of our atomistic model. Note that for each of these diffraction patterns, we conducted parallel simulations on 50 independent configurations to enhance statistical reliability.
III.2 Dynamic properties
III.2.1 Structural relaxation
In this section, we explore the SISF and the structural relaxation or $\alpha$-relaxation time of the dodecagonal quasicrystal. Fig. 10(a) illustrates the temporal evolution of CR-$F_{s}(k^{*},t)$ as the temperature increases, calculated using Eq. (10). Here, the SISF is defined at the wave vector $k=7.3$, which corresponds to the major peak in $S(k)$ and is sensitive to variations in atomic density, making it an important indicator of structural transitions, particularly in disordered phases. SISF focuses on the timescale for the decay of density-density correlations of matter, and it is typically defined at the first peak of $S(k)$. The evolution of the SISF reveals a two-step relaxation process, including the short-time $\beta$ relaxation and the long-time $\alpha$-relaxation. The $\alpha$-relaxation time $\tau_{\alpha}$ is defined as the time when $F_{s}(k^{*},\tau_{\alpha})=1/e$, as indicated by the horizontal dashed line in Fig. 10(a).
It is evident that structural relaxation, as measured by the self-intermediate scattering function, occurs in our quasicrystalline material, is similar to both glass-forming liquids [91] and heated crystalline solids [85] in the sense that there is a general fast beta relaxation occurring on a timescale on the order of a ps that is relatively temperature insensitive and a longer time alpha relaxation process having a relaxation time that is highly temperature dependent.
As the temperature decreases, the $\alpha$-relaxation time increases rapidly, a well-known, even defining, attribute to glass-forming liquid, but this behavior is also characteristic to heated crystals and derives from the reduced kinetic of the atoms in the crystal and the associated reduced rate of atomic diffusion [85, 85]. Eventually, at a certain crossover temperature in the liquid regime (not study in the present paper), the $\alpha$- and fast $\beta$-relaxation processes merge, defining an onset temperature $T_{A}$ for the multi-step dynamics exhibited by glass-forming, crystalline and quasi-crystal materials. This pattern of relaxation thus seem to be a universal pattern of relaxation in condensed matter rather broadly. At low temperatures, the dynamics of the quasicrystal becomes so slow that it becomes increasingly difficult to make simulation observations under equilibrium conditions, a familiar situation in glass-forming liquids, but again this is the case in crystalline materials [92]. To quantity the observations in Fig. 10, a necessary step in comparing to the dynamics of crystalline and glass-forming materials sufficiently to enable determination which material is more like our quasi-crystalline system, we fit our $F_{s}(k,t)$ data to a two-step relaxation function, defined as,
This function has been found to describe $F_{s}\left(k^{*},t\right)$ for a range of materials [93, 94].
Our fitting of the fast dynamics properties $\tau_{f}$ and $\beta_{f}$ leads to $\tau_{f}\approx 0.1$ and $\beta_{f}\approx 1.9$ (Fig. 10(b)), which is rather typical of glass-forming liquids [95]. The ”mode amplitude” or ”non-ergodicity” parameter $A_{\alpha}$ displays relatively weak temperature dependence within the explored temperature range. Importantly, fixing these three parameters ($\tau_{f}$, $\beta_{f}$, and $A_{\alpha}$) does not significantly compromise the accuracy of our fits across the entire temperature range investigated[94], allowing for precise quantification of our quasi-crystal relaxation. At this stage, it is not evidently clear whether our quasi-crystal material is more like a ”typical” crystalline material or a glass-forming liquid as these materials share many features in their dynamics when temperature is varied. We next consider the temperature dependence of $\tau_{\alpha}$ to search for some discriminating characteristics of the quasi-crystal dynamics that could help us in this determination.
In Fig. 10(c), we present the $T$ dependence of $\tau_{\alpha}$ in our 2D dodecagonal quasicrystal in the traditional form of an Arrhenius plot. We observe a relatively high temperature regime in which $\log\tau_{\alpha}$ exhibits roughly linear variation with $1/T$, but the slope bends over sharply as the $T$ decreases below 0.5, a temperature notably in the quasicrystal phase regime. We emphasis here that we did not simulate our quasi-crystalline material in the liquid regime so that should be another Arrhenius-like regime for $\tau_{\alpha}$ at higher temperature, with probably a different slope (activation energy). We thus simulate an intermediate temperature range where multi-step relaxation regime and cannot assign the onset temperature of glass-forming corresponding to a temperature $T_{A}$ where the $\alpha$ and fast $\beta$ relaxation processes start to merge, and where relaxation can be reliably be taken as being Arrhenius and non-cooperative to a good approximation [96, 97]. Qualitatively, the kink-like transition from Arrhenius to non-Arrhenius behavior aligns with numerous previous observations on some glass-forming materials [98, 99], but also highly anharmonic crystalline materials, such as superionic crystalline materials [100]. (We will discuss this Arrhenius curve ”kink” further below in relation to its genreal occurrence in the intermediate forms of matter mentioned by Angell [10].)Thus, we still have not arrived at the desired descrimination of quasi-crystalline materials in relation to other more studied condensed materials.
Past experience on both glass-forming liquids [95] and heated crystalline materials [100] has shown the estimation of the activation energy from the local slope of this type of Arrhenius plot cannot be reliably used to define the activation energy under conditions in which the slope of the Arrhenius curve varies with $T$ can be highly misleading [95]. An estimate of the activation free energy for diffusion requires that the actual $T$ dependent activation free energy $\Delta G(T)$, formally defined by the relation, $\Delta G(T)\equiv k_{B}T\ln(\tau_{\alpha}/\tau_{0})$ [97, 101, 95]. Here, $\tau_{0}$ serves as a prefactor associated with the vibrational relaxation time, estimated at approximately 0.05 in our reduced units. Estimates of the temperature dependent activation energy deduced from this definition are indicated in the inset of Fig. 10(c) where we observe that $\Delta G(T)$ progressively increases upon cooling over the temperature range that we investigate. We interpret this trend to be consistent with the sigmoidal temperature variation of $\Delta G(T)$ observed in previous studies of structural relaxation of Al-Sm, a glass-forming liquid [97, 95] that exhibits fragile to strong type glass-formation, and experiments and simulation on superionic crystalline materials [100]. Again, we have not been able to resolve whether the dynamics is more like a crystal or a heated crystalline system, forcing us the nature of dynamic heterogeneity in the quasi-crystal material in the hopes of finally resolving the situation.
III.2.2 Dynamic heterogeneity
To capture the “dynamic heterogeneity”, i.e., mobility fluctuations of the dodecagonal quasicrystal, we employ the non-Gaussian parameter frequently used in the literature, $\alpha_{2}\left(t\right)$ [102, 103]. This parameter quantifies the degree of deviation of particle displacements from a Gaussian distribution and this phenomenon has often been found in association with large mobility fluctuations in complex liquids, making this quantity an indirect measure of such dynamic heterogeneity. The time evolution of CR-$\alpha_{2}\left(t\right)$ for the dodecagonal quasicrystal is depicted in Fig. 11(a) at various temperatures. Similar to glass-forming liquids, $\alpha_{2}(t)$ exhibits a peak at a characteristic time $t^{*}$ [102] that is generally shorter than $\tau_{\alpha}$. This measure of the ”lifetime” of the dynamic heterogeneity increases upon decreasing temperature, a trend typical of glass-forming liquids, but contrary to heated crystals where this quantity appears to increase upon heating from the low temperature crystal state where this quantity is very small, but then passes through a maximum at intermediate temperatures and then falls upon approaching the melting temperature[85]. This decreases of $\alpha_{2}$ upon cooling was found to be especially prevalent in the simulation of superionic ${\rm UO}_{2}$ in its high temperature [100]. This is the first hint that the dynamics of quasi-crystalline material is perhaps more like a glass-forming liquid at low $T$ than a heated crystal. However, we do not consider the $T$ regime at which the hexatic phase of the quasicrystalline material enters the liquid states, the analog of the melting in the 3-dimensional crystalline material studied previously. In contrast, we find that $\alpha_{2,\text{max}}$ progressively increases upon lowering $T$, a pattern of behavior that aligns with the dynamics of glass-forming liquids rather than crystalline materials [104]. This trend clearly points to a commonality between glass-forming liquids and our quasi-crystalline material. We next consider the interrelation between diffusion and the structural relaxation time and $t^{*}$, which as we shall see offers further and more compelling evidence for a closer relationship between the dynamics of our quasi-crystalline material and glass-forming material rather than the dynamics of crystalline materials.
III.2.3 Revisiting the kink in the Arrhenius curve of relaxation
In Fig. 12(a), we show the evolution of the temperature dependence of CR-MSD for the dodecagonal quasicrystal when the temperature is varied over a large range. As in condensed material rather generally, particle MSD exhibits three stages that mirrors these same dynamical regimes in $F_{s}(k^{*},t)$. At short times, particles undergo ballistic motions, i.e., a particle does not collide with any particles at very short times so that the motion resembles particles in a gas. During this stage, the distance traveled varies nearly linearly with time, resulting in MSD being proportional to $t^{2}$. Beyond this ”inertial dynamics” regime, a plateau in $\left\langle\Delta r^{2}\left(t\right)\right\rangle$ develops which becomes more pronounced at lower temperatures. This pattern of behavior is characteristic of condensed materials rather broadly so it is no surprise we see the same basic trend in our quasi-crystal material. At long times, the particles escape from their ”cages” at intermediate time in which most of the atoms vibrate within regions defined by surrounding particles. The Debye-Waller parameter $\left\langle u^{2}\right\rangle$, which describes the amplitude of atomic vibrations on the fast $\beta$ relaxation time (see Fig. 10(a)), defining the onset of caging and the ”end” of the fast $\beta$ relaxation process. At sufficiently long times, particle motion can be described as simple diffusion, provided the material is in thermal equilibrium. Of course, the timescale required for this long time diffusion emerges, corresponding to a time on the order of $\tau_{\alpha}$ can be extremely large and difficult to access by molecular dynamics simulation. This can make the estimation of the diffusion derived from particle displacement difficult to estimate computationally as we shall next show.
While this method of estimating $D$ provides a reasonable method for estimating $D$ in our quasi-crystalline material, the slowing of the dynamics limits the temperature range in which estimates can be made and we confine our estimates to the same $T$ range as we considered above for $\tau_{\alpha}$. Fig. 12(b) shows the result of our fitting procedure based on Eq. (16) to estimate $D$ as function of $T$.
As in the case of $\tau_{\alpha}$, we see that $D$ exhibits two distinct temperature regimes when the data is considered in the form of an Arrhenius plot, as depicted in Fig. 12(c). Measurements of $D$ in a real quasi-crystalline material have indicated a striking similar ”kink” in the Arrhenius curve for $D$ [50], confirming this interesting feature found in our simulations. Recent simulations of $D$ for a model metallic glass material [95] also strikingly resemble our observations in Fig. 12, where a kink change arises in the Arrhenius curve for $D$ data indicating a clear change in the dynamical regimes of the material where the origin of this change is investigated in depth [95, 101]. Arrhenius curves having a ”kink” as in Fig. 12 having upward curvature characteristically arise in the dynamics of enzymes [106, 107, 108], an observation in line with Angell’s suggestion that proteins could also be viewed as an intermediate form of matter having properties in common with other materials of this ypte. It was mentioned before that this type of ”kink” in the Arrhenius curves is found in Al-Sm metallic glass-forming liquids [95, 101] and moreover this type of non-Arrhenius dynamics is observed in simulations and measurements of the dynamics of model superionic materials, and the interfacial dynamics of crystalline materials [109], providing further evidence supporting Angell heuristic arguments.
Below we show that the temperature variation of $D$ and $\tau_{\alpha}$ bares a direct relation to $t^{*}$ and $\tau_{\alpha}$ so we do not provide a discussion of the $T$ dependent activation energy as it must be the same as deduced from Fig. 12(c). However, we do point out, contrary to the visual impression of the Arrhenius plots of Fig. 10(c) and Fig. 12(c), that the activation energy monotonically increases upon cooling rather than decreasing upon cooling.
Figure 13: Characterization of the atomic motion mechanisms in quasicrystal by displacement vector (yellow arrows) at different temperatures. The emergence of collective 1D string-like structures is evident at high temperatures.
III.2.4 Atomic-scale mechanisms governing relaxation and diffusion in the fast $\beta$ and $\alpha$ relaxation regimes
To further investigate the mechanisms of the transition between a distinct fast $\beta$ dynamics regime and $\alpha$relaxation and the curious kinks in the Arrhenius curves for $\tau_{\alpha}$ and $D$, based on a direct examination on the shape of particle trajectories and the occurrence of collective motion in which more than one particle is involved. First, we examine the atomic trajectories and correlated motion at different temperatures over $3\times 10^{7}$ MD steps. At the lowest $T$ considered, the most prevalently observed atomic displacements, apart from relatively small-scale vibrations, involve isolated local particle hopping, that might be compatible with hypothetical “phason flips” suggested to exist in quasicrystal systems [110]. These ”defects” transform energetically comparable configurations into others and remind us of moving interstitials in crystalline materials. Larger scale collective motions might possibly at these low temperatures, but if they do occur, they must be rare. Starting about $T=0.5$, string-like particle exchange motion or ”string-like” motion emerges. In particular, this form of collective motion involving a permutational exchange motion starts becomes conspicuous. Previous simulations on both crystalline materials and glass-forming liquids exhibit what appears to be a superficially similar collective motion so this type of motion to be rather general. Importantly, the $T$ regimes associated with these distinct regimes of particle diffusion and relaxation align with the kink observed in Fig. 10(c) and Fig. 12(c).
We next investigate the particle displacement process in greater detail. In Figure 14, we present an illustration of a putative phason flip at the very low temperature, $T=0.3$. Notably, all observed jumps occur at structural imperfections within the quasicrystal. In the case of a hexagonal tile, nonadjacent particles exist in a relatively unstable state, as predicted by the LJG potential. Consequently, particles apparently tend to move in search of more energetically favored positions, providing an energetic impetus for the phason flips. Figures 14(a) and 14(b) depict the spatial distribution of the same particle cluster at $t=0$ and $t=3\times 10^{4}$, respectively. The red atom undergoes a phason flip (or several reversible phason flips) during this period. Interestingly, regardless of how many times these flips occur, the form of defects, i.e., the hexagonal tiles, persists and no topological transition is observed. Moreover, the number of defects exhibits a strong correlation with the preparation process of the quasicrystal phase, specifically the cooling rate. Quasicrystals with lower potential energy, corresponding to a lower concentration of defects, can be obtained through a slower cooling rate. In practice, forming ”perfect” quasicrystals without defects is challenging due to the limitations of the timescale and length scale of our atomistic simulations, making it difficult to reach an equilibrium concentration of defects.
@@ III.2.3 Revisiting the kink in the Arrhenius curve of relaxation
Figure 14: Illustration of the phason flip mechanism at low temperature $T=0.3$. (a) and (b) are the spatial distribution of a particle cluster at $t=0$ and $t=3\times 10^{4}$, respectively. The yellow arrow represents the displacement vector of the red particle which has been scaled 1.5 times for clear observation.
[…]
@@
Figure 16: Minimum energy pathway of a phason flip. The yellow arrows represent displacement vectors which are scaled twice for clear observation.
@@
Figure 17: Minimum energy pathway of a string-like motion in form of a single loop. The yellow arrows represent the real displacement vectors.
@@
Figure 18: Minimum energy pathway of a string-like motion in form of a ruddged line. The yellow arrows represent the real displacement vectors.
III.2.5 Energy landscape of atomic motion
@@
First, Figure 16 illustrates the MEP of what we temtatively identify as a phason flip. The displacement process apparently involves an atom climbing over a single energy saddle point. We select specific characteristic points along the MEP to observe the detailed atomic displacement form. As the energy increases, the atom gradually moves into the position of an existing defect. Upon reaching the saddle point F, a new form of defect appears. Subsequently, the atom gradually settles into a new equilibrium position. The activation energy required for this entire process is approximately 2.5, aligning well with the data obtained from the Arrhenius law in Fig. 12 at the low-temperature regime.
@@
In the case of string-like motion, depicted in Figs. 17 and 18, the energetic landscape traversed evidently involves a much greater complexity, the MEP involves multiple saddle points and several meta-stable positions. The illustrative ring-like exchange motion initiates with the movements of two or three particles of the hexagonal ring, accompanied by the generation of a defect left in the tail (Fig. 17D). This forms the first meta-stable state, with $\Delta E$ for this process being about 6, nearly three times the value for a phason flip. Subsequently, it evolves into a semi-loop form (Fig. 17E) with the disappearance of the produced defect in Fig. 17C. This cluster tends to be also meta-stable, causing the MEP to drop to a point F. At this point, the structure and bonds differ slightly from the initial configuration after traveling half the distance of the nearest neighbor separation, with the energy of state F being higher than that in the initial state. No further meta-stable state is encountered after this point in the collective motion evolution, apparently prompting the cluster of atoms to continue its motion in a collective way. Ultimately, a full loop is completed, and the structure returns to its initial equilibrium position.

### v2

@@ Acknowledgement
Heat Capacity
We calculate the heat capacity using the “direct” method of taking the derivative of the total energy with respect to temperature in cooling paths. Although this method, while applicable in non-equilibrium heating/cooling paths, may large uncertainties in estimating thermodynamic properties in equilibrium [111]. It is sufficient for calculating the relative value of different phases and representing phase transitions. In a manner similar to the density, the heat capacity as a function of temperature (Fig. 13) allows us to precisely estimate the location of the thermodynamic transitions of the quasicrystal material. These transition temperatures ($T_{LH}=1.78$ and $T_{HQ}=0.87$) are consistent with previous findings. While both of these transitions seem sharp, it is not possible to deduce the transition order due to the limited system size, as mentioned previously.
@@
Radial distribution function
The RDF is a widely utilized tool for elucidating the structural characteristics of ordered and disordered materials. In the case of crystalline materials, $g(r)$ exhibits distinct peaks corresponding to the periodic arrangements of particles, signifying the presence of long-range order. For isotropic systems, $g(r)$ relies solely on the magnitude of the relative atomic distances and can be calculated as follows:
$g\left(r\right)=\frac{1}{2\pi\,r\,\Delta r}\frac{1}{\rho}\left\langle\sum_{i\neq i^{\prime}}\delta\left(r-\lvert\mathbf{r}_{i}-\mathbf{r}_{i^{\prime}}\rvert\right)\right\rangle,$ (A1)
where $\rho$ represents the number density.
In Fig. 14, we present the RDF for the liquid phase at $T=1.95$, the hexatic phase at $T=1.2$, and the dodecagonal quasicrystal at $T=0.3$. The RDF of the quasicrystal exhibits a distinctive behavior that positions it between the characteristics of crystalline and disordered materials. This observation supports the notion of an intermediate state of matter, as alluded to in the title and introduction of our paper.
In addition to the well-established RDF profiles in amorphous and hexatic phases, the quasicrystal exhibits unique features. In the short range, the RDF of the quasicrystal reveals distinct peaks, which can be attributed to the quasi-periodic tiling. The first peak emerges at approximately $r=1.0$, while the second peak appears around $r=1.89$. These values are readily explicable as they correspond to the positions of potential energy wells (stable states) within the LJG-type potential [refer to Fig. 1(a)]. On the other hand, at longer distances the RDF of the quasicrystal displays typical characteristics of disordered materials. It gradually converges towards a constant value with non-vanishing fluctuations, reflecting the absence of long-range translational symmetry, akin to amorphous systems, while still retaining a certain translational order.
@@
Static structure factor
@@
Figure 15: Static structure factor of the (a) liquid phase at $T=1.95$, (b) hexatic phase at $T=1.2$ and (c) dodecagonal quasicrystal at $T=0.3$. The rapid oscillations in (b) and (c) arise from insufficient statistics of small $k$ data with finite simulation size.
As depicted in Fig. 15, the static structure factors of the liquid, hexatic, and quasicrystal phases all prominently feature a main peak around $k^{*}=2\pi/a$, where $a$ represents the nearest-neighbor distance in real space. A pronounced peak in the structure factor at $k^{*}$ serves as a indication of local correlations in the particle positions in the condensed material. This peak corresponds to a wave vector that generates a long-range oscillatory behavior with a periodicity of $a$ in real space [113]. Moreover, the height of the first peak in the static structure factor is a critical parameter for predicting the liquid-solid phase transition. According to the Hansen-Verlet criterion in crystalline materials the freezing transition occurs when the value of the first peak in $S(k)$ reaches a specific constant value [114]. In the liquid phase, the first peak is notably lower than those in the hexatic and quasicrystal phases, underscoring both the lack of long-range order and the inherent disorder in the system [115].
Appendix B: Collective motion in dodecagonal quasicrystal
@@
Figure 16: Illustration of defect hopping at low temperature $T=0.3$. (a) and (b) are the spatial distribution of a particle cluster at $t=0$ and $t=3\times 10^{4}$, respectively. The yellow arrow represents the displacement vector of the red particle which has been scaled 1.5 times for clear observation. A supplementary video 1 is provided to visualize the dynamic process of such defect hopping.
We investigate the particle displacement process during fast $\beta$ relaxation at lower temperatures. In Fig. 16, we present an illustration of a putative defect flip at the very low temperature, $T=0.3$. Notably, all observed jumps occur at structural imperfections within the quasicrystal. In the case of a hexagonal tile, nonadjacent particles exist in a relatively unstable state, as predicted by the LJG potential. Consequently, particles apparently tend to move in search of more energetically favored positions, providing a possible energetic impetus for defect flips. Fig. 16(a) and 16(b) depict the spatial distribution of the same particle cluster at $t=0$ and $t=3\times 10^{4}$, respectively. The red atom undergoes a “jump” during this period. Interestingly, regardless of how many times these putative defect flips occur, the form of defects, i.e., the hexagonal tiles, persists and no topological transition is observed. The whole dynamic process is shown in Supplemental Video 2. Moreover, the number of defects exhibits a strong correlation with the preparation process of the quasicrystal phase, specifically the cooling rate. Quasicrystals with lower potential energy, corresponding to a lower concentration of defects, can be obtained through a slower cooling rate. In practice, forming “perfect” quasicrystals without defects is challenging due to the limitations of the timescale and length scale of our atomistic simulations, making it difficult to reach an equilibrium concentration of defects.
Appendix C: Energy landscape of atomic motion
@@
Figure 17: Minimum energy pathway of a defect flip. The yellow arrows represent displacement vectors which are scaled twice for clear observation. The atomistic configurations from ’A’ to ’H’ shown in the right panels are corresponding to positions of the same letters indicated on the energy pathway in the left panel.
@@
Figure 18: Minimum energy pathway of a string-like motion in form of a single loop. The yellow arrows represent the real displacement vectors. The atomistic configurations from ’A’ to ’J’ shown in the right panels are corresponding to positions of the same letters indicated on the energy pathway in the left panel.
@@
Figure 19: Minimum energy pathway of a string-like motion in form of a crooked line. The yellow arrows represent the real displacement vectors. The atomistic configurations from ’A’ to ’L’ shown in the right panels are corresponding to positions of the same letters indicated on the energy pathway in the left panel.
@@
First, Fig. 17 illustrates the MEP of what we tentatively identify as a phason flip. The displacement process apparently involves an atom climbing over a single energy saddle point. We select specific characteristic points along the MEP to observe the detailed atomic displacement form. As the energy increases, atoms gradually move into the position of an existing defect. Upon reaching the saddle point F, a new form of defect appears. Subsequently, the atom gradually settles into a new equilibrium position. The activation energy required for this entire process is approximately 2.5, aligning well with the data obtained from the Arrhenius law in Fig. 7 at the low-temperature regime.
@@
In the case of string-like motion, depicted in Figs. 18 and 19, the energetic landscape traversed evidently involves a much greater complexity, the MEP involves multiple saddle points and several meta-stable positions. The illustrative ring-like exchange motion initiates with the movements of two or three particles of the hexagonal ring, accompanied by the generation of a defect left in the tail (Fig. 18D). This leads to the formation of an initial meta-stable state, with $\Delta E$ for this process being about 6, nearly three times the value assigned to the putative phason flip. Subsequently, it evolves into a semi-loop form (Fig. 18E) with the disappearance of the produced defect in Fig. 18C. This cluster tends to be also meta-stable, causing the MEP to drop to a point F. At this point, the structure and bonds differ slightly from the initial configuration after traveling half the distance of the nearest neighbor separation, with the energy of state F being higher than that in the initial state. No further meta-stable state is encountered after this stage in the collective motion evolution, apparently prompting the cluster of atoms to continue its motion in a collective way. Ultimately, a full loop is “completed”, and the structure returns to its initial equilibrium position.